1. (10) Solve the following initial-value problem. Give an implicit solution (x + ye /*)dx - xe /*dy = 0, y(1) = 0
Answer #1

solution :

Given. (*+ years) de-xeht, dy zo ; 90-0 e dy = (x + y et) dx = x el =) atyyl dy xe पात्र dx 7 dy -9/ е yle dy (xté ५) da LetInty= x 0 (05 (4)+c) = 4:09 (0:) -). 9(*): plog ( 00g) +C) =): 0 :-) ICS! :-) Y (X) = x log (log(68) +.c) =) Y(x) = x log (lo